a 2-digit number that can be rounded to 80 whose some of the digits is 14

Answer by KMST(5330)

The numbers 75 through 84 would be rounded to 80.
Of those, the only one whose sum of digits is 14 is 77, because 7+7=14.
As you increase the number by 1 from 75 to 84,
the sums of digits increase by 1 from 75 to 79,
for 75, 7+5=12,
for 76, 7+6=13,
for 77, 7+7=14,
for 78, 7+8=15,
for 79, 7+9=16,
and then the sum falls down to 8 for 80, and starts increasing by 1 again,
for 80, 8+0=8,
for 81, 8+1=9,
for 82, 8+2=10,
for 83, 8+3=11, and
for 84, 8+4=12.